# Estimation of relative risk, odds ratio and their logarithms with guaranteed accuracy and controlled sample size ratio

### v2

@@ 4.2 Group sampling
Thus, proceeding as in Section 2.2.2 and taking into account (141),
@@
where $\operatorname{E}[G]\bar{\phi}$ is given by (153). Figure 18 represents the theoretical approximation (155), as well as results from simulation. As with element sampling, the efficiency for large $A$ is slightly lower than in RR and LRR (Figures 11 and 14), and for small $A$ there is almost no difference between simulation results and theoretical approximation even if $\phi$ is large.
@@ 5.1 Element sampling
The estimation method for the LOR, $\Psi=\log(p_{1}(1-p_{2})/(p_{2}(1-p_{1})))$, has few differences compared to that presented for OR in Section 4. The second stage estimates $\log(p_{1}/(1-p_{1}))$ and $\log(p_{2}/(1-p_{2}))$ separately, using for each the method described in Mendo [16, section 3], which consists of two IBS procedures with the same parameter. Specifically, to estimate $\log(p_{i}/(1-p_{i}))$, $i=1,2$, IBS is applied to population $i$ until $s_{i}$ successes are obtained, which requires a random number $N_{i}^{\prime}$ of observations from this population, and then IBS is applied until $s_{i}$ failures are obtained, which requires $N_{i}^{\prime\prime}$ additional observations. These are the same steps as for OR, but with $\alpha=0$. Then, conditioned on $s_{1}$ and $s_{2}$, which are obtained in the first stage, $H_{N_{i}^{\prime}-1}+H_{N_{i}^{\prime\prime}-1}$ is a conditionally unbiased estimator of $\log(p_{i}/(1-p_{i}))$ with conditional variance less than $1/(s_{i}-5/4)$ for any $p_{i}\in(0,1)$ [16, theorem 3]. Therefore,
@@
$\operatorname{Var}[\hat{\Psi}^{2}\mid s_{1},s_{2}]<\frac{1}{s_{1}-5/4}+\frac{1}{s_{2}-5/4}.$ (157)
@@
The LOR estimator uses $\xi=1$ and $r_{1}$ given by (81), which in this case, taking into account (159), is written as
@@
As with previous estimators, this choice of $r_{1}$ yields average numbers of samples close to their minimum values with respect to this parameter (results not shown). The curve $c(A,r_{1},1)=0$, plotted in Figure 3, is almost indistinguishable from that for OR; and (160) corresponds to rounding up the ordinate values of this curve, with a minimum of $3$.
@@
Particularizing (87) for $\zeta=\Psi$, with $\partial\Psi/\partial p_{1}=1/\bar{p}_{1}$, $\partial\Psi/\partial p_{2}=-1/\bar{p}_{2}$, and then using (148), (149) and the fact that $\operatorname{Var}[\hat{\Psi}]<A$, it follows that the efficiency with element sampling is bounded by (151) (with the value of $\mu_{1}$ corresponding to LOR and with $r_{1}$ computed accordingly). Figure 19 shows the results, which are very similar to those for OR.
@@ 5.2 Group sampling
The efficiency with group sampling is expressed by (155), as for OR. The results, plotted in Figure 20, are not the same as in that case because of the differences in $\mu_{1}$ and $r_{1}$, which affect $\operatorname{E}[G]$ and thus the efficiency; but they are seen to be very similar.
@@ 6 Conclusions
As an example, consider the estimation of $p_{1}p_{2}$. A straightforward approach would be to generate samples with parameter $p_{1}p_{2}$ from samples with parameters $p_{1}$ and $p_{2}$, using a procedure analogous to the Bernoulli factory described in Section 4, and then apply IBS to the generated samples. However, this does not give any control on the proportion of samples from the two populations, and the average number of samples required from each population scales with the inverse of $p_{1}p_{2}$. Instead, the method presented for RR and LRR can be used: two IBS processes with parameters $s_{1}$ and $s_{2}$ are respectively applied to the two populations (second stage), which requires $N_{1}$ and $N_{2}$ samples, where $s_{1}$ and $s_{2}$ are obtained from a previous pair of IBS processes with fixed parameters $r_{1}$ and $r_{2}$ (first stage). Then,
@@ Appendix A Estimation procedure and properties of the estimators
The estimation procedure for RR and LRR, and that for OR and LOR, are respectively described in Algorithms 1 and 2, which also provide a summary of their properties.
@@
    Compute $r_{1}$ from (81) with $c(A,r_{1},\xi)$ given by (76) (for LRR the explicit expression (119) can equivalently be used), or obtain it by rounding up the value from Figure 3, with a minimum of $3$. Compute $r_{2}$ as $r_{1}+\mu_{1}-\mu_{2}$.
@@
    Compute $r_{1}$ from (81) with $c(A,r_{1},\xi)$ given by (145) (for LOR the explicit expression (160) can equivalently be used), or obtain it by rounding up the value from Figure 3, with a minimum of $3$. Set $r_{2}=r_{1}$.
@@
    (Second sampling stage): From population $1$, take as many samples as needed, $N^{\prime}_{1}$, to obtain $s_{1}$ successes; then as many as needed, $N^{\prime\prime}_{1}$, to obtain $s_{1}-\alpha$} failures. From population $2$, take as many samples as needed, $N^{\prime}_{2}$}, to obtain $s_{2}-\alpha$} successes; then as many as needed, $N^{\prime\prime}_{2}$, to obtain $s_{2}$ failures. Let $N_{1}=N_{1}^{\prime}+N_{1}^{\prime\prime}$, $N_{2}=N_{2}^{\prime}+N_{2}^{\prime\prime}$.
@@ B.1 For relative risk and log relative risk
$\begin{split}\lim_{\phi\rightarrow 0}\Pr[M_{1}\phi\leq x]&=\lim_{\phi\rightarrow 0}\sum_{k=r_{1}}^{\lfloor x/\phi\rfloor}\binom{k-1}{r_{1}-1}\left(\phi\sqrt{\theta}\right)^{r_{1}}\left(1-\phi\sqrt{\theta}\right)^{k-r_{1}}\\ &=\frac{\theta^{r_{1}/2}}{(r_{1}-1)!}\int_{0}^{x}t^{r_{1}-1}\exp(-t\sqrt{\theta})\,\mathrm{d}t.\end{split}$ (161)
@@
$\operatorname{Cov}\left[X,-\frac{1}{X}\right]=-1+\operatorname{E}\left[X\right]\operatorname{E}\left[\frac{1}{X}\right]\approx\frac{2r_{1}+\mu_{1}-\mu_{2}-1}{(r_{1}-1)(r_{1}+\mu_{1}-\mu_{2}-1)}.$ (170)
@@
It will be established next that (166), (168) and the right-hand side of (172) are much smaller than (171). To this end, the following observations will be useful. From (60) and (63) it stems that $r_{1}\approx r_{2}=r_{1}+\mu_{1}-\mu_{2}$ with little approximation error. In practice, the target $A$ will typically be smaller, or much smaller, than $1$. For example, a relative RMSE of $10\%$ corresponds to $A=0.01$. On the other hand, $\xi=1$ is small compared to $1/A$, and so are the values of $r_{1}$ obtained from (81); namely $r_{1}\approx\sqrt{1/A}$, as can be seen from (82) or in Figure 3. In the example, $A=0.01$ gives $r_{1}=9$ for RR and $10$ for LRR.
To show that (166) is much smaller than (171), it is convenient to study the cases $\lambda\theta\approx 1$, $\lambda\theta\gg 1$ and $\lambda\theta\ll 1$ separately. For $\lambda\theta\approx 1$, the above approximations imply that (166) and (171) reduce to $2r_{1}$ and $8(1/A+r_{1})^{2}/r_{1}$ respectively, and their ratio,
@@
which is again small compared with $1$. For $\lambda\theta\ll 1$ the result is the same. Thus (168) can also be neglected to a good approximation.
@@
The conclusion of the preceding analysis is that (165) can be approximated by keeping only the third summand in the right-hand side, as it is significantly larger than the others. This means that the variability in $\Delta$ is mostly due to the variability of $N_{1}$ and $N_{2}$ caused by variations of $s_{1}$ and $s_{2}$, i.e. (iii) as defined above. The variability of $M_{1}$ and $M_{2}$, (i), and the variability of $N_{1}$ and $N_{2}$ conditioned on $s_{1}$ and $s_{2}$, (ii), are comparatively smaller. Therefore, to compute $\operatorname{E}[|\Delta|]$, the variables $M_{1}$, $M_{2}$ in (92) can be replaced by their means, and $N_{1}$, $N_{2}$ can be replaced by their conditional means given $s_{1}$, $s_{2}$, which yields (94).
@@ B.2 For odds ratio and log odds ratio
The variables $N_{1}^{\prime}$, $N_{1}^{\prime\prime}$, $N_{2}^{\prime}$, $N_{2}^{\prime\prime}$, and therefore $N_{1}$, $N_{2}$, depend on $M_{1}$, $M_{2}$ only through $\bar{M}_{1}$, $\bar{M}_{2}$. This implies that $M_{1}$, $M_{2}$, $N_{1}$, $N_{2}$, are conditionally independent given $\bar{M}_{1}$, $\bar{M}_{2}$. Thus, using (92), the first term in (173) is written as
@@
This is similar to the decomposition (165) in RR and LRR, where only the covariance term is different. The approximate expressions of the summands for small $\phi$ are also different from that case, as will be computed next. Then, as in RR an LRR, it will be shown that for OR and LOR the third summand dominates the other three.
@@
The first stage applies IBS to samples with parameter $\bar{p}_{i}$, using a number $\bar{M}_{i}$ of those samples, of which $r_{i}$ are successes. Each sample with parameter $\bar{p}_{i}$ is generated by the Bernoulli factory described in Section 4.1, taking samples with parameter $p_{i}$ as inputs. With this factory, the average number of inputs needed to produce an output is $3/2$; a success output always uses $2$ inputs, and a failure output uses either $1$ or $2$ inputs. Let $\pi_{i}$ denote the probability that $2$ inputs are used, conditioned on the output being a failure. Then $3/2=2\bar{p}_{i}+(1+\pi_{i})(1-\bar{p}_{i})$, from which
@@
The second summand in (177) is computed next. Using the fact that $N_{1}^{\prime}$, $N_{1}^{\prime\prime}$, $N_{2}^{\prime}$, $N_{2}^{\prime\prime}$ are conditionally independent given $s_{1}$, $s_{2}$, taking into account (13) and approximating for $\phi$ small,
$\begin{split}&\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\\ &\ =\frac{\operatorname{E}[s_{1}](1-p_{1})\bar{\phi}^{2}}{\lambda p_{1}^{2}}+\frac{\operatorname{E}[s_{1}-\alpha]p_{1}\bar{\phi}^{2}}{\lambda(1-p_{1})^{2}}+\frac{\lambda\operatorname{E}[s_{2}-\alpha](1-p_{2})\bar{\phi}^{2}}{p_{2}^{2}}+\frac{\lambda\operatorname{E}[s_{2}]p_{2}\bar{\phi}^{2}}{(1-p_{2})^{2}}\\ &\ =\left(\frac{\operatorname{E}[s_{1}]((1-p_{1})^{3}+p_{1}^{3})-\alpha p_{1}^{3}}{\lambda\bar{p}_{1}^{2}}+\frac{\lambda\left(\operatorname{E}[s_{2}]((1-p_{2})^{3}+p_{2}^{3})-\alpha(1-p_{2})^{3}\right)}{\bar{p}_{2}^{2}}\right)\bar{\phi}^{2}\\ &\ =\frac{\operatorname{E}[s_{1}](1-3\bar{p}_{1})-\alpha p_{1}^{3}}{\lambda\bar{\theta}}+\lambda\bar{\theta}\left(\operatorname{E}[s_{2}](1-3\bar{p}_{2})-\alpha(1-p_{2})^{3}\right)\\ &\ \approx\frac{\operatorname{E}[s_{1}]}{\lambda\bar{\theta}}+\lambda\bar{\theta}(\operatorname{E}[s_{2}]-\alpha)\end{split}$ (186)
@@
Regarding the third summand, from (11) it stems that $\operatorname{E}[N_{1}]=\operatorname{E}[N_{1}^{\prime}]+\operatorname{E}[N_{1}^{\prime\prime}]=s_{1}/\bar{p}_{1}-\alpha/(1-p_{1})$, and similarly $\operatorname{E}[N_{2}]=s_{2}/\bar{p}_{2}-\alpha/p_{2}$. Then, using (37) and (38),
@@
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\\ \approx\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)^{\!2}(2r_{1}-1)\cdot\left(\frac{1}{r_{1}(r_{1}-2)}\left(\lambda\bar{\theta}+\frac{1}{\lambda\bar{\theta}}\right)+\frac{2}{r_{1}^{2}}\right).$ (189)
The fourth summand in (177) is bounded using the Cauchy-Schwarz inequality as in Appendix B.1, together with (181) and the fact that $(3-4\bar{p}_{i})/(2(1-\bar{p}_{i}))$ is less than $3/2$ (and approaches that value for $\phi$ small):
@@
It is concluded from the above analysis that the variability of $\Delta$ in OR and LOR, as in RR and LRR, is mostly caused by that of $s_{1}$ and $s_{2}$. This implies that to compute $\operatorname{E}[|\Delta|]$ each variable $M_{i}$, $i=1,2$ in (92) can be replaced by its mean, given by (132), and $N_{i}$ can be replaced by its conditional mean given $s_{i}$, which is approximately $s_{i}/\bar{p}_{i}$ according to (120) and (121). This establishes (152).

### v3

@@ 4.2 Group sampling
Thus, proceeding as in Sect. 2.2.2 and taking into account (141),
@@
where $\operatorname{E}[G]\bar{\phi}$ is given by (153). Figure 18 represents the theoretical approximation (155), as well as results from simulation. As with element sampling, the efficiency for large $A$ is slightly lower than in RR and LRR (Figs. 11 and 14), and for small $A$ there is almost no difference between simulation results and theoretical approximation even if $\phi$ is large.
@@ 5.1 Element sampling
The estimation method for the LOR, $\Psi=\log(p_{1}(1-p_{2})/(p_{2}(1-p_{1})))$, has few differences compared to that presented for OR in Sect. 4. The second stage estimates $\log(p_{1}/(1-p_{1}))$ and $\log(p_{2}/(1-p_{2}))$ separately, using for each the method described in Mendo [16, section 3], which consists of two IBS procedures with the same parameter. Specifically, to estimate $\log(p_{i}/(1-p_{i}))$, $i=1,2$, IBS is applied to population $i$ until $s_{i}$ successes are obtained, which requires a random number $N_{i}^{\prime}$ of observations from this population, and then IBS is applied until $s_{i}$ failures are obtained, which requires $N_{i}^{\prime\prime}$ additional observations. These are the same steps as for OR, but with $\alpha=0$. Then, conditioned on $s_{1}$ and $s_{2}$, which are obtained in the first stage, $H_{N_{i}^{\prime}-1}+H_{N_{i}^{\prime\prime}-1}$ is a conditionally unbiased estimator of $\log(p_{i}/(1-p_{i}))$ with conditional variance less than $1/(s_{i}-5/4)$ for any $p_{i}\in(0,1)$ [16, theorem 3]. Therefore,
@@
$\operatorname{Var}[\hat{\Psi}\mid s_{1},s_{2}]<\frac{1}{s_{1}-5/4}+\frac{1}{s_{2}-5/4}.$ (157)
@@
The LOR estimator, like previous ones, uses $\xi=1$ and $r_{1}$ given by (81), which in this case, taking into account (159), is written as
@@
As with previous estimators, this choice of $r_{1}$ yields average numbers of samples close to their minimum values with respect to this parameter (results not shown). The curve $c(A,r_{1},1)=0$, plotted in Fig. 3, is almost indistinguishable from that for OR; and (160) corresponds to rounding up the ordinate values of this curve, with a minimum of $3$.
@@
Particularizing (87) for $\zeta=\Psi$, with $\partial\Psi/\partial p_{1}=1/\bar{p}_{1}$, $\partial\Psi/\partial p_{2}=-1/\bar{p}_{2}$, and then using (148), (149) and the fact that $\operatorname{Var}[\hat{\Psi}]<A$, it follows that the efficiency with element sampling is bounded by (151) (with the value of $\mu_{1}$ corresponding to LOR and with $r_{1}$ computed accordingly). Figure 19 shows the results, which are very similar to those for OR (Fig. 17).
@@ 5.2 Group sampling
The efficiency with group sampling is expressed by (155), as for OR. The results, plotted in Fig. 20, are not the same as in that case (Fig. 18) because of the differences in $\mu_{1}$ and $r_{1}$, which affect $\operatorname{E}[G]$ and thus the efficiency; but they are seen to be very similar.
@@ 6 Conclusions
As an example, consider the estimation of $p_{1}p_{2}$. A straightforward approach would be to generate samples with parameter $p_{1}p_{2}$ from samples with parameters $p_{1}$ and $p_{2}$, using a procedure analogous to the Bernoulli factory described in Sect. 4, and then apply IBS to the generated samples. However, this does not give any control on the proportion of samples from the two populations, and the average number of samples required from each population scales with the inverse of $p_{1}p_{2}$. Instead, the method presented for RR and LRR can be used: two IBS processes with parameters $s_{1}$ and $s_{2}$ are respectively applied to the two populations (second stage), which requires $N_{1}$ and $N_{2}$ samples, where $s_{1}$ and $s_{2}$ are obtained from a previous pair of IBS processes with fixed parameters $r_{1}$ and $r_{2}$ (first stage). Then,
@@ Appendix A Estimation procedure and properties of the estimators
The estimation procedure for RR and LRR, and that for OR and LOR, are respectively described in Algorithms 1 and 2. The properties of the estimators are also summarized.
@@
    Compute $r_{1}$ from (81) with $c(A,r_{1},\xi)$ given by (76) (for LRR the explicit expression (119) can equivalently be used), or obtain it by rounding up the value from Fig. 3, with a minimum of $3$. Compute $r_{2}$ as $r_{1}+\mu_{1}-\mu_{2}$.
@@
    Compute $r_{1}$ from (81) with $c(A,r_{1},\xi)$ given by (145) (for LOR the explicit expression (160) can equivalently be used), or obtain it by rounding up the value from Fig. 3, with a minimum of $3$. Set $r_{2}=r_{1}$.
@@
    (Second sampling stage): From population $1$, take as many samples as needed, $N^{\prime}_{1}$, to obtain $s_{1}$ successes; then as many as needed, $N^{\prime\prime}_{1}$, to obtain $s_{1}-\alpha$} failures. From population $2$, take as many samples as needed, $N^{\prime}_{2}$}, to obtain $s_{2}-\alpha$} successes; then as many as needed, $N^{\prime\prime}_{2}$, to obtain $s_{2}$ failures. Let $N_{1}=N_{1}^{\prime}+N_{1}^{\prime\prime}$ and $N_{2}=N_{2}^{\prime}+N_{2}^{\prime\prime}$.
@@ B.1 For relative risk and log relative risk
$\begin{split}\lim_{\phi\rightarrow 0}\Pr[M_{1}\phi\leq x]&=\lim_{\phi\rightarrow 0}\sum_{k=r_{1}}^{\lfloor x/\phi\rfloor}\binom{k-1}{r_{1}-1}\left(\phi\sqrt{\theta}\right)^{r_{1}}\left(1-\phi\sqrt{\theta}\right)^{k-r_{1}}\\ &=\frac{\theta^{r_{1}/2}}{(r_{1}-1)!}\int_{0}^{x}t^{r_{1}-1}\exp(-t\sqrt{\theta})\,\mathrm{d}\mkern 0.8mut.\end{split}$ (161)
@@
$\operatorname{Cov}\left[X,-\frac{1}{X}\right]=-1+\operatorname{E}\left[X\right]\operatorname{E}\left[\frac{1}{X}\right]\approx\frac{r_{1}+r_{2}-1}{(r_{1}-1)(r_{2}-1)}.$ (170)
@@
It will be established next that (166), (168) and the right-hand side of (172) are much smaller than (171). To this end, the following observations will be useful. From (60) and (63) it stems that $r_{1}\approx r_{2}=r_{1}+\mu_{1}-\mu_{2}$ with little approximation error. In practice, the target $A$ will typically be smaller, or much smaller, than $1$. For example, a relative RMSE of $10\%$ corresponds to $A=0.01$. On the other hand, $\xi=1$ is small compared to $1/A$, and so are the values of $r_{1}$ obtained from (81); namely $r_{1}\approx\sqrt{1/A}$, as can be seen from (82) or in Fig. 3. In the example, $A=0.01$ gives $r_{1}=9$ for RR and $10$ for LRR.
To show that (166) is much smaller than (171), it is convenient to study the cases $\lambda\theta\approx 1$, $\lambda\theta\gg 1$ and $\lambda\theta\ll 1$ separately. For $\lambda\theta\approx 1$, the approximations in the above paragraph imply that (166) and (171) reduce to $2r_{1}$ and $8(1/A+r_{1})^{2}/r_{1}$ respectively, and their ratio,
@@
which is again small compared with $1$. For $\lambda\theta\ll 1$ the result is the same. Thus the second summand in (165) can also, to a good approximation, be neglected in comparison with the third.
@@
The conclusion of the preceding analysis is that (165) can be approximated by keeping only the third summand in the right-hand side, as it is significantly larger than the others. This means that the variability in $\Delta$ is mostly due to the variability of $N_{1}$ and $N_{2}$ caused by variations of $s_{1}$ and $s_{2}$, i.e. (iii) as defined above. The variability of $M_{1}$ and $M_{2}$, (i), and that of $N_{1}$ and $N_{2}$ conditioned on $s_{1}$ and $s_{2}$, (ii), are comparatively smaller. Therefore, to compute $\operatorname{E}[|\Delta|]$, the variables $M_{1}$, $M_{2}$ in (92) can be replaced by their means, and $N_{1}$, $N_{2}$ can be replaced by their conditional means given $s_{1}$, $s_{2}$, which yields (94).
@@ B.2 For odds ratio and log odds ratio
The variables $N_{1}^{\prime}$, $N_{1}^{\prime\prime}$, $N_{2}^{\prime}$, $N_{2}^{\prime\prime}$, and therefore $N_{1}$, $N_{2}$, depend on $M_{1}$, $M_{2}$ only through $\bar{M}_{1}$, $\bar{M}_{2}$. This implies that $M_{1}$, $M_{2}$, $N_{1}$, $N_{2}$, are conditionally independent given $\bar{M}_{1}$, $\bar{M}_{2}$. Thus, using (92), the first term in the right-hand side of (173) is written as
@@
This is similar to the decomposition (165) in RR and LRR, where only the covariance term is different. Although the approximate expressions of the summands for small $\phi$, to be computed next, are different from that case, it will be shown that for OR and LOR the third summand again dominates the other three.
@@
The first stage applies IBS to samples with parameter $\bar{p}_{i}$, using a number $\bar{M}_{i}$ of those samples, of which $r_{i}$ are successes. Each sample with parameter $\bar{p}_{i}$ is generated by the Bernoulli factory described in Sect. 4.1, taking samples with parameter $p_{i}$ as inputs. With this factory, the average number of inputs needed to produce an output is $3/2$; a success output always uses $2$ inputs, and a failure output uses either $1$ or $2$ inputs. Let $\pi_{i}$ denote the probability that $2$ inputs are used, conditioned on the output being a failure. Then $3/2=2\bar{p}_{i}+(1+\pi_{i})(1-\bar{p}_{i})$, from which
@@
The second summand in (177) is computed as follows. Using the fact that $N_{1}^{\prime}$, $N_{1}^{\prime\prime}$, $N_{2}^{\prime}$, $N_{2}^{\prime\prime}$ are conditionally independent given $s_{1}$, $s_{2}$, taking into account (13) and approximating for $\phi$ small,
$\begin{split}&\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\\ &\ =\frac{\operatorname{E}[s_{1}](1-p_{1})\bar{\phi}^{2}}{\lambda p_{1}^{2}}+\frac{\operatorname{E}[s_{1}-\alpha]p_{1}\bar{\phi}^{2}}{\lambda(1-p_{1})^{2}}+\frac{\lambda\operatorname{E}[s_{2}-\alpha](1-p_{2})\bar{\phi}^{2}}{p_{2}^{2}}+\frac{\lambda\operatorname{E}[s_{2}]p_{2}\bar{\phi}^{2}}{(1-p_{2})^{2}}\\ &\ =\left(\frac{\operatorname{E}[s_{1}]((1-p_{1})^{3}+p_{1}^{3})-\alpha p_{1}^{3}}{\lambda\bar{p}_{1}^{2}}+\frac{\lambda\left(\operatorname{E}[s_{2}]((1-p_{2})^{3}+p_{2}^{3})-\alpha(1-p_{2})^{3}\right)}{\bar{p}_{2}^{2}}\right)\bar{\phi}^{2}\\ &\ =\frac{\operatorname{E}[s_{1}](1-3\bar{p}_{1})-\alpha p_{1}^{3}}{\lambda\bar{\theta}}+\lambda\bar{\theta}\left(\operatorname{E}[s_{2}](1-3\bar{p}_{2})-\alpha(1-p_{2})^{3}\right)\\ &\ \approx\frac{\operatorname{E}[s_{1}]}{\lambda\bar{\theta}}+\lambda\bar{\theta}(\operatorname{E}[s_{2}]-\alpha),\end{split}$ (186)
@@
As for the third summand, using (120)–(122) and then taking into account (37) and (38),
@@
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\\ \approx\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)^{\!2}(2r_{1}-1)\left(\frac{1}{r_{1}(r_{1}-2)}\left(\lambda\bar{\theta}+\frac{1}{\lambda\bar{\theta}}\right)+\frac{2}{r_{1}^{2}}\right).$ (189)
The fourth summand in (177) is bounded by means of the Cauchy-Schwarz inequality as in Appendix B.1; and then, using (181) and the fact that $(3-4\bar{p}_{i})/(2(1-\bar{p}_{i}))$ is less than $3/2$ (and approaches that value for $\phi$ small),
@@
It is concluded from the above analysis that the variability of $\Delta$ in OR and LOR, as in RR and LRR, is mostly caused by that of $N_{1}$ and $N_{2}$ induced by the variations of $s_{1}$ and $s_{2}$. This implies that to compute $\operatorname{E}[|\Delta|]$ each variable $M_{i}$, $i=1,2$ in (92) can be replaced by its mean, given by (132), and $N_{i}$ can be replaced by its conditional mean given $s_{i}$, which is approximately $s_{i}/\bar{p}_{i}$ according to (120) and (121). This establishes (152).
Acknowledgments
The author would like to thank an anonymous reviewer for helpful comments, the editor-in-chief Prof. Ulrike Schneider for her efficient handling of the review process, and Dr. Ben O’Neill for advice on terminology.